OpenStudy (anonymous):

Your weight lifting class cost you a $10 fee up front and $5 per class after. Y = 5x + 10. What is the y-intercept for in the equation?

OpenStudy (anonymous):

so this is already in y=mx+b form where b is the y intercept

OpenStudy (anonymous):

so what is b in the equation you typed?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

to find the y-intercept simply set x to be 0 in your eqaution and solve for b (since you have the form y=mx+b) Y = 5x + 10 y=5(0)+10 y=0+10 y=10